# Thermodynamic and observational constraints on black holes with primary hair in Beyond Horndeski Gravity: Stability and shadows

### v1

@@ Abstract
We found that, unlike General Relativity, the shift-symmetric subclass of Beyond Horndeski theories provides black holes with primary hair, which are thermodynamically stable and compatible with current Event Horizon Telescope observations for M87* and Sgr A* black holes. We first investigate its thermodynamic properties by studying the influence of the primary hair on the thermodynamic quantities and the local stability, imposing severe constraints in the allowed range of the primary hair values. Then, we analyze the null geodesics near this black hole, demonstrating how the scalar hair influences the shadow diameter, finding that higher values of the primary hair lead to smaller shadow sizes. Additionally, we constrain the scalar hair using the observational data, revealing its sensitivity to the choice of other black hole parameters. To further probe the observational features of the scalar hair, we simulate face-on two-dimensional images of spherically infalling accretion disks, examining how the primary scalar hair affects the black hole’s shadow. Finally, we impose all relevant constraints to identify black holes that are both stable and consistent with observational data.
@@ II Brief review on the black hole solution with primary scalar hair
In the framework of Beyond Horndeski theories [56], where the additional shift symmetry $\varphi\rightarrow\varphi+\mathrm{const.}$ and the parity symmetry $\varphi\rightarrow{-\varphi}$ are present, the theories are parametrized by three arbitrary functions $G_{2}$, $G_{4}$ and $F_{4}$, which are given in terms of the kinetic term of the scalar field $X=-\frac{1}{2}\varphi_{,\mu}\varphi^{,\mu}$, and hence, the action is given by
$I\left[g_{\mu\nu},\varphi\right]=\frac{1}{2}\int\mathrm{d}^{4}x\,\sqrt{-g}\Biggl{\{}G_{2}(X)+G_{4}(x)R+G_{4X}\left[\bigl{(}{\varphi^{;\mu}}_{;\mu}\bigr{)}^{2}+\varphi_{,\mu;\nu}\varphi^{,\mu;\nu}\right]+F_{4}(X)\,\epsilon^{\mu\nu\rho\sigma}{\epsilon^{\alpha\beta\gamma}}_{\sigma}\varphi_{,\mu}\varphi_{,\alpha}\varphi_{,\beta;\nu}\varphi_{,\gamma;\rho}\Biggr{\}},$ (1)
where $G_{4X}\equiv\partial_{X}G_{4}$. In Ref. [47], they found that the Horndeski functionals that support the hairy black hole are given by,
@@ IV Null geodesics around a black hole with primary scalar hair and the shadow constraints
The diagrams indicate that as the scalar hair increases, the photon sphere shrinks, affecting the final shadow size observed in the sky. Additionally, larger values of the ratio $M/\lambda$ correspond to larger photon spheres. For $\mathcal{c}=0$ (the SBH), the well-known values are $r_{p}=3M$ for the photon sphere radius and $b_{c}=3\sqrt{3}M$ for the critical impact parameter.
@@
The absence of a lower theoretical bound for Sgr A* shadow diameter begins at a specific value of $M/\lambda$, which correlates with the changes in the photon sphere radius, as shown in Fig. 7(b). Beyond this value, the lowest possible theoretical shadow diameter is higher than the lower bound constrained by $1\sigma$ confidence intervals for Sgr A*. This critical value of $M/\lambda$ is highlighted in Fig. 10(a), where the range of acceptable values for $\mathcal{c}$ and $M/\lambda$ that align with the EHT observations is depicted. After this critical value, the upper bound for $\mathcal{c}$ compatible with Sgr A* observations is governed by $c_{*}=4M/(\pi\lambda)$.
@@ VI Summary and conclusions
This paper has studied a black hole spacetime characterized by primary scalar hair, focusing on its thermodynamical properties, null geodesics, and observational signatures. Our findings reveal the significant influence that scalar hair can exert on the physical characteristics of s and their observable features. We began by examining the thermodynamic behavior of the black hole, where we observed that the presence of primary scalar hair introduces notable modifications to the black hole’s temperature, entropy, and specific heat. In particular, the primary hair generates intermediate-size black hole branches that are locally stable and not close to Schwarzschild black in the parameters space. Namely, we have found that locally stable black holes exist only for
@@
In our analysis of null geodesics, we discovered that scalar hair significantly influences the shadow cast by the black hole. The shadow size decreases as we increase the value of $\mathcal{c}=\lambda q^{4}$. This means that in a given theory with $\lambda$ fixed, those black holes with higher primary hair develop wider shadows. This observation is particularly relevant given the recent imaging data from the EHT. By applying the EHT constraints to our model, we found that the scalar hair parameter can be bounded within certain limits, depending on the black hole’s mass-to-coupling ratio. Our results demonstrate that the shadow size is highly sensitive to both the magnitude and sign of the scalar hair and other black hole parameters. This sensitivity underscores the importance of accurately determining these parameters to constrain the presence of scalar hair through observational data effectively. Remarkably, the family of theories given by $\lambda$ and $eta$ always contain black holes with primary hair in agreement with both EHT observed shadows, and such range on the primary hair increases as the mass of the black hole is higher.
@@
Our study provides compelling evidence that primary scalar hair profoundly impacts the thermodynamic and observational properties of black holes. Constraints from local thermodynamic stability and constraints obtained by observations can be intersected. Doing so, we found that, unlike in General Relativity, the primary hair allows locally thermodynamically stable black holes whose shadow sizes are compatible with current EHT observations. The sensitivity of black hole shadow and accretion disk images to scalar hair, as constrained by EHT data, opens up new possibilities for testing the existence of such hair through direct astronomical observations. As the resolution and sensitivity of black hole imaging techniques continue to advance, the influence of scalar hair could become a detectable feature, offering a unique probe into the nature of black holes and the fundamental theories that describe them. More research is needed to extend these analyses to rotating black holes and explore how the theory’s parameters affect circularity. Additionally, further observational efforts should aim to refine the constraints on scalar hair using more extensive datasets from EHT and other upcoming interferometric arrays. These endeavors hold the potential to significantly enhance our ability to test alternative theories of gravity and deepen our understanding of black hole physics.

### v2

@@ Abstract
In this paper, we find that unlike in General Relativity, the shift-symmetric subclass of Beyond Horndeski theories permits black holes with primary hair that are thermodynamically stable and align with current Event Horizon Telescope observations of the M87* and Sgr A* black holes. This work begins by investigating thermodynamic properties, analyzing how primary hair influences thermodynamic quantities and local stability, which imposes strict constraints on the allowed range of primary hair values. The null geodesics near this black hole are then examined, demonstrating how scalar hair affects the shadow diameter. Specifically, when the parameter of the Beyond Horndeski function $F_{4}$ is negative, increasing scalar hair enlarges the shadow; in contrast, when this parameter is positive, greater scalar hair reduces the shadow size. Further constraints on the scalar hair are derived using observational data, highlighting its sensitivity to other black hole parameters. To explore additional observational features, face-on two-dimensional images of spherically infalling accretion disks are simulated, revealing how primary scalar hair shapes the black hole’s shadow. Finally, all relevant constraints are combined to identify black holes that are both stable and consistent with observational data.
@@ II Brief review on the black hole solution with primary scalar hair
This section provides an overview of the key features of the static, spherically symmetric black hole solution found in Ref. [47]. In the framework of Beyond Horndeski theories [56], where the additional shift symmetry $\varphi\rightarrow\varphi+\mathrm{const.}$ and the parity symmetry $\varphi\rightarrow{-\varphi}$ are present, the theories are parametrized by three arbitrary functions $G_{2}$, $G_{4}$ and $F_{4}$, which are given in terms of the kinetic term of the scalar field $X=-\frac{1}{2}\varphi_{,\mu}\varphi^{,\mu}$. This action is given by
$I\left[g_{\mu\nu},\varphi\right]=\frac{1}{2}\int\mathrm{d}^{4}x\,\sqrt{-g}\Biggl{\{}G_{2}(X)+G_{4}(X)R+G_{4X}\left[(\square\varphi)^{2}-\varphi_{\mu\nu}\varphi^{\mu\nu}\right]+F_{4}(X)\epsilon^{\mu\nu\rho\sigma}\epsilon^{\alpha\beta\gamma}{}_{\sigma}\varphi_{\mu}\varphi_{\alpha}\varphi_{\nu\beta}\varphi_{\rho\gamma}\Biggr{\}},$ (1)
where the following notations are adopted for conciseness: $\varphi_{\mu}=\partial_{\mu}\varphi$, $\varphi_{\mu\nu}=\nabla_{\mu}\partial_{\nu}\varphi$ and derivation with respect to $X$ is represented by a subscript $X$. The Horndeski functionals that support the hairy black hole are given by,
@@ IV Null geodesics around a black hole with primary scalar hair and the shadow constraints
The diagrams indicate that as the $\mathcal{c}$-parameter increases, the photon sphere shrinks, affecting the final shadow size observed in the sky. Additionally, larger values of the ratio $M/\lambda$ correspond to larger photon spheres. For $\mathcal{c}=0$ (the SBH), the well-known values are $r_{p}=3M$ for the photon sphere radius and $b_{c}=3\sqrt{3}M$ for the critical impact parameter.
@@
The absence of a lower theoretical bound for Sgr A* shadow diameter begins at a specific value of $M/\lambda$, which correlates with the changes in the photon sphere radius, as shown in Fig. 7(b). Beyond this value, the lowest possible theoretical shadow diameter is higher than the lower bound constrained by $1\sigma$ confidence intervals for Sgr A*. This critical value of $M/\lambda$ is highlighted in Fig. 10(a), where the range of acceptable values for $\mathcal{c}$ and $M/\lambda$ that align with the EHT observations is depicted. After this critical value, the upper bound for $\mathcal{c}$ compatible with Sgr A* observations is governed by $\mathcal{c}_{*}=4M/(\pi\lambda)$.
@@ VI Summary and conclusions
In this article, we explored a black hole spacetime endowed with primary scalar hair, examining its thermodynamic properties, null geodesics, and observational signatures. Our results revealed a profound influence of scalar hair on the black hole’s physical properties, particularly its shadow and observable characteristics. We began by examining the thermodynamic behavior of the black hole, where we observed that the presence of primary scalar hair introduces notable modifications to the black hole’s temperature, entropy, and specific heat. In particular, the primary hair generates intermediate-size black hole branches that are locally stable and not close to Schwarzschild black in the parameters space. Namely, we have found that locally stable black holes exist only for
@@
In our analysis of null geodesics, we discovered that scalar hair significantly influences the shadow cast by the black hole. We find that the shadow size decreases as the parameter $\mathcal{c}=\eta q^{4}$ is increased from negative to positive values. Thus, the effect of primary hair on the shadow size, within a given theory with fixed $\eta$, can be explained as follows: in the negative domain of $\mathcal{c}$, where $\eta<0$, a larger scalar hair leads to a larger shadow. Conversely, in the positive domain of $\mathcal{c}$, where $\eta>0$, a larger scalar hair results in a smaller shadow (see Fig. 8). This observation is particularly relevant given the recent imaging data from the EHT. By applying the EHT constraints to our model, we found that the scalar hair parameter can be bounded within certain limits, depending on the black hole’s mass-to-coupling ratio. Our results demonstrate that the shadow size is highly sensitive to both the magnitude and sign of the scalar hair and other black hole parameters. This sensitivity underscores the importance of accurately determining these parameters to effectively constrain the presence of scalar hair through observational data. Remarkably, the family of theories given by $\lambda$ and $\eta$ always contain black holes with primary hair in agreement with both EHT observed shadows, and such range on the primary hair increases as the mass of the black hole is higher.
@@
Our study provides compelling evidence that primary scalar hair profoundly impacts the thermodynamic and observational properties of black holes. Constraints from local thermodynamic stability and constraints obtained by observations can be intersected. In doing so, we found that, unlike general relativity, the primary hair allows locally thermodynamically stable black holes whose shadow sizes are compatible with current EHT observations. The sensitivity of black hole shadow and accretion disk images to scalar hair, as constrained by EHT data, opens up new possibilities for testing the existence of such hair through direct astronomical observations. As the resolution and sensitivity of black hole imaging techniques continue to advance, the influence of scalar hair could become a detectable feature, offering a unique probe into the nature of black holes and the fundamental theories that describe them.
Certainly, more research is needed to extend these analyses to rotating black holes and explore how the theory’s parameters affect circularity. Additionally, further observational efforts should aim to refine the constraints on scalar hair using more extensive datasets from EHT and other upcoming interferometric arrays. It is important to note that, in this work, we have used the black hole shadow size as the main tool to constrain the black hole parameters. As a result, no specific constraint has been placed on the parameter $\lambda$. To refine the observational constraints further, it is also essential to consider additional experimental tests, such as the period and periapsis shift in the orbits of S-stars near the center of the Milky Way. Incorporating such new experimental data could provide a more comprehensive framework for testing the black hole model and constraining its parameters. These endeavors hold the potential to significantly enhance our ability to test alternative theories of gravity and deepen our understanding of black hole physics.